# Dynamic Neural Curiosity Enhances Learning Flexibility for Autonomous Goal Discovery

### v1

@@ Abstract
The autonomous learning of new goals in robotics remains a complex issue to address. Here, we propose a model where curiosity influence learning flexibility. To do so, this paper proposes to root curiosity and attention together by taking inspiration from the Locus Coeruleus-Norepinephrine system along with various cognitive processes such as cognitive persistence and visual habituation. We apply our approach by experimenting with a simulated robotic arm on a set of objects with varying difficulty. The robot first discovers new goals via bottom-up attention through motor babbling with an inhibition of return mechanism, then engage to the learning of goals due to neural activity arising within the curiosity mechanism. The architecture is modelled with dynamic neural fields and the learning of goals such as pushing the objects in diverse directions is supported by the use of forward and inverse models implemented by multi-layer perceptrons. The adoption of dynamic neural fields to model curiosity, habituation and persistence allows the robot to demonstrate various learning trajectories depending on the object. In addition, the approach exhibits interesting properties regarding the learning of similar goals as well as the continuous switch between exploration and exploitation.
Index Terms:
Intrinsic motivation, curiosity, attention, task engagement, locus coeruleus, autonomous goal learning, dynamic neural fields.
@@ I Introduction
Developmental robotics takes inspiration from various fields such as developmental psychology, neuroscience, machine learning or philosophy. A popular domain of interest in the field and one of the core idea of this study is called Curiosity. Curiosity can be seen as a particular case of Intrinsic Motivation (IM) and researchers have demonstrated that IM could be an efficient drive for the autonomous learning of behavior [1]. In addition, the learning of various behavior exhibits a developmental pattern [2, 3]. Nowadays, there are multiple ways of modelling curiosity and one of them consists of training predictors such as forward and inverse models and thus displaying a learning progress that depends on the error variation.
@@
Along with curiosity, attention is a fundamental cognitive process. Indeed, attention allows humans to focus and concentrate on specific elements of our surroundings. However, attention is a broad term gathering several specific mechanisms. For example, visual attention is the ability to sustain the gaze at salient stimuli and is developing during the first months after birth [4], [5]. To help search for specific visual stimuli, inhibition of return is a well studied mechanism naturally preventing a person to look twice at the same location within a short interval of time [6]. In addition, we can distinguish two types of visual attention : exogenous, driven by external stimuli or endogenous and goal-driven [7], [8]. These two types of attention are often referred as bottom-up for the former and top-down for the later, and inhibition of return can happen during both, but whether there are shared neural pathways between them in order to occur is still subject to debate [9]. In the brain, arousal is mediated by the Locus Coeruleus-Norepinephrine (LC-NE) and demonstrates different patterns of neural activity. According to the adaptive gain theory [10], LC neurons exhibit two modes of activation : tonic and phasic. During tonic activation, the subject disengages from the current task and engages in an exploratory behavior. On the contrary, phasic LC activation engages in exploitation and thus is driven by task specific outcomes. It has been demonstrated that IM is involved in tonic and phasic activation of dopamine [11], overlapping certain neural pathways with noradrenaline (norepinephrine) [12] and thus might be directly interacting with the LC-NE system. If curiosity and more especially its learning progress component is related to attention, then the LC-NE system is a potential starting point.
@@
Here, we propose a robotic cognitive architecture for the online discovery and learning of goals where a robotic arm is learning how to interact with simple objects. To do so, we introduce a model situating the LC-NE system at the center of different processes, thus determining if the robot should explore and discover potential goals through bottom-up attention, or exploit and learn to achieve these goals. During both exploration and exploitation, an inhibition of return mechanism generates a set of actions toward the object location until producing a change in the environment. If the inhibition of return is shared by exploration and exploitation, there is however a conceptual difference between them since the discovery of a new goal in the environment rests on exogenous attention, and on a top-down and goal driven approach for the learning of these goals. For the exploration stage, we complete the attentional mechanism connected to the LC-NE system with an habituation paradigm [13], where the discovery of a new goal produces a tonic activation and limits the time of exploration. To balance this process, the absence of a new goal signals a shift from tonic to phasic activation. During the latter, the learning of a goal occurs with the use of simple neural networks implementing a forward and inverse model. The output from these neural networks determine an error associated with a goal and its learning progress. With the exception of the forward and inverse models, the entire architecture is designed with dynamic neural fields [14]. Goals are then represented within neural fields by the neural activation of their learning progress and are projected to the LC-NE system, which then chooses to exploit the goal with the highest learning progress. By design, the model follows an enactive approach by continuously pushing the robot to interact with its environment and thus shaping its cognition by using several forms of working memory. We will demonstrate that when close goals are discovered, the robot autonomously only focuses on the one with the highest error and slowly forgets about the other. We evaluate several neural parameters such as persistence, inhibition of error and habituation that modulate and optimize the learning of goals, depending on the objects. To our knowledge, this is the first attempt to model intrinsic motivation with dynamic neural fields and thus support the robot to express different learning behaviors by tuning intrinsic properties of the dynamic neural fields.
@@
This paper will provide a theoretical background according to the literature in Section II before presenting the architecture of the experiments (Section III). Then, we will present the experimental setup with the frameworks used in Section IV followed by the different results (Section V). Finally, we will discuss about the advantage and limitations of the architecture (Section VI) before concluding.
@@ II-A Curiosity and Intrinsic motivation
In terms of biological perspective, intrinsic motivation (IM) needs to be distinguished from extrinsic motivation [15]. The former can be seen as a way for the brain to monitor the learning efficiency of a particular knowledge while the latter directly guides the learning itself through external cues/rewards. Then, there are different metrics to compute intrinsic motivation [16], separating knowledge-based from competence-based systems [17]. There is a consensus in the literature classifying curiosity as a particular case of intrinsic motivation [18]. Indeed, curiosity focuses on the learning progress hypothesis [19], [1] and formulates it as the evolution of prediction error while learning a task. In practice, curiosity can be used for the learning of a world model by leveraging and focusing the sensorimotor experience [20]. GRAIL [21] is a cognitive architecture operating on different levels where the IM used is competence-based, determining the best goal to pursue depending on how well the learning occurs. Intrinsic motivation in robotics is also used to form hypotheses in developmental psychology. For example, a developmental robotics model identified IM to explain the sudden tool-use improvement in children [22]. However, the discovery of goals in visually rich environments is problematic due to the exploration space becoming quickly too vast for an intrinsically motivated system to be efficient. To address this issue, researchers are applying a set of dimension reduction methods. One of the first study to overcome these limitations is the case of a robot arm interacting with a ball [23]. In this case, a variational autoencoder is trained and goals are sampled from the resulting goal space. Other work in the domain and related to robot grasping extended this concept by linking and compressing the object features, the action and the sensori outcomes together [24]. By doing so, the robot creates a latent space where the learning of forward and inverse models is facilitated to achieve grasping. Treating with high dimensional spaces with intrinsic motivation is far from trivial, and the sampling problem can gain insights with a more developmental point of view. In these studies, it is necessary to perform several round of exploration in order to build up the different goal learning progress and then select them for learning. However, it is not clear how the switch between discovery and learning takes place in infancy. In our work, we propose that the error contributes at some extend to ease this shift. Finally, IM proved to be a drive reproducing developmental patterns [25], but actively and efficiently sampling the environment remain difficult to tackle [26]. However, researchers demonstrated that IM involves other cognitive processes, such as attention, and provide perspective toward possible neural basis of intrinsic motivation [27].
@@ II-B Attention and inhibition of return
Visual attention is essential for an individual to focus on an unexpected event and to optimize cognitive processes on a specific task [8]. This distinction separates exogenous (bottom-up) from endogenous (top-down) attention. If motivation can sharpen exogenous spatial attention under certain conditions [28], the modulation of top-down attention can enhance the learning of a particular task [29, 30, 31]. The neural basis for these distinct processes are, for a significant part, shared [32], but not entirely [33]. Regarding the coupling between attention and intrinsic motivation, a study used reinforcement learning to allocate and shift attention [34]. More close to our approach, researchers used bottom-up attention to learn object affordances and decompose a task into sub-goals [35]. Here, we choose to focus on the inhibition of return effect (IOR) to facilitate attention toward an object and ease the discovery of new goals. In the literature, the general mechanism of exploration is often resumed as random and direct exploration [36, 37] where the former consists of producing random actions with a high uncertainty but potentially large expected reward. The latter directs exploration toward more certain rewards but necessitates prior information. With curiosity in robotics, this difference can be translated to a motor babbling behavior for random exploration and the building of representations for direct exploration. For example, a robot can learn a representation of a diverse set of goals encountered during random exploration to generate new goals [24]. Here, our purpose is to propose an additional exploration method that can reduce the uncertainty of random exploration via the inhibition of return effect. This mechanism was pointed out by Posner [6] in visual attention when he discovered that the return of attention to a cue previously attended expresses a longer time. On the opposite, there is a facilitation (faster response time) if the time interval between cue-targets is short. This consequence endows the brain with a natural foraging mechanism of visual stimuli [38], but it remains necessary to proceed to an attentional disengagement in order to observe the IOR effect [39]. However, it appears that an endogenous disengagement of attention is not enough in order to observe the effect and is instead always relying on exogenous disengagement of attention [40]. In addition, it is possible to induce an inhibitory of return for both exogenous (uncued location) [41], [9] and endogenous (cued location) attention, but the effect seems highly correlated to the cue’s saliency and thus relies on an important exogenous component [42]. Regarding IOR in action production, there is no consensus whether the effect is directly or partially involved. Several research reproduced reaction time of IOR during a hand reaching movement [43] and even showed that reaching paths were biased toward the cue locations under certain conditions [44]. Directly connected to our research, it has been determined that IOR effects can be applied to objects [45]. Even more, inhibition can interact with working memory to mediate goal-directed action toward an object, optimizing performance for future interaction. In consequence, we intend to generate and use IOR around the object location so the robot can generate goal-driven actions. To be able to identify these two forms of attention and decide to engage in exploration, we take inspiration from the Locus Coeruleus.
II-C Exploration, exploitation and the Locus Coeruleus system
@@
The locus coeruleus-norepinephrine system (LC-NE) is a brainstem nuclei that is widely connected to the neocortex. The first unifying approach regarding its role and functions is the adaptive gain theory [10]. The theory demonstrates two different neural activations, depending on the arousal state as well as the relation with task engagement. More precisely, the LC-NE system exhibits a tonic neural activation during exploration, rising arousal’s level and thus bottom-up attention toward salient events [46]. Here, the exploration strategy consists of generating goal poses around the object with IOR and observe the possible outcomes. A phasic activation suggests a goal-oriented behavior (exploitation) with a focus on the task (top-down attention). Then, the LC-NE system is a crossroad between attentional processes and decision making. Aston-Jones and Cohen determine the emergence of a tonic activation when the utility of a task vanishes. The LC-NE receives direct connection from the anterior cingulate (Acc) and orbitofrontal cortices (OFC) which are directly involved in task-related utility and decision making [47], [48]. More specifically, the ACC seems to represent value predictions about reward and, more generally, uncertainty [49]. There is a large consensus to recognize the basal ganglia as the processing place of rewards that can support reinforcement learning [50]. Several robotics experiments take inspiration from the phasic delivery of dopamine to signal rewards based on prediction errors [51], [52]. We can observe a similarity between the phasic delivery of noradrenaline from the LC and the phasic dopamine-based signal emerging from the basal ganglia, especially after pointing out that dopamine and noradrenaline share neural pathways [53]. In addition, researchers demonstrate that if both dopamine and noradrenaline are related to rewards prediction, only the latter is predictive of task engagement and is strongly activated when cues indicate a new task condition. This means that on the contrary to dopamine, noradrenaline is strongly involved in motivating current or future engagement. The Acc directly projects prediction information about the task utility to the LC-NE system which in turn decides if the utility is high enough to drive the neural activation in a phasic mode, hence to an exploitatory stage. Following the predictive coding approach from Friston [54], researchers proposed a model of active inference base on the LC-NE system’s ability to track prediction errors by linking states and actions together [55]. The interesting point of the study is how the LC can possibly rely on a prediction error mechanism and thus be related to intrinsic motivation. Under the prism of active inference, they demonstrate that the two firing modes of the LC are an emergent property from the processing of state-action prediction errors. Furthermore, scientists used functional magnetic resonance imaging (fMRI) to determine the neural substrates of intrinsic motivation [56]. They suggest that diminished patterns of neural activity correlate with individuals exhibiting high intrinsic motivation. This finding does not specify which brain region is concerned by intrinsic motivation, yet high intrinsic motivation seems to rely on a phasic neural activation. Therefore, we believe that the functioning of the LC system is promising when using curiosity for robotics. It will allow the robot to engage in exploration and exploitation depending on uncertainty, meaning on how accurate the robot’s predictions are, based on the learning progress. As a consequence, the LC model will follow a tonic or phasic neural activation, thus autonomously tuning attention to modulate the level of arousal and engage in exploratory activity or the learning of a skill.

### v2

@@ Abstract
The autonomous learning of new goals in robotics remains a complex issue to address. Here, we propose a model where curiosity influences learning flexibility. To do so, this paper proposes to root curiosity and attention together, taking inspiration from the Locus Coeruleus-Norepinephrine system along with various cognitive processes such as cognitive persistence and visual habituation. We apply our approach by experimenting with a simulated robotic arm on a set of objects with varying difficulty. The robot first discovers new goals via bottom-up attention through motor babbling with an inhibition of return mechanism, then engages in the learning of goals due to neural activity arising within the curiosity mechanism. The architecture is modeled with dynamic neural fields, and the learning of goals such as pushing objects in diverse directions is supported by the use of forward and inverse models implemented by multilayer perceptrons. The adoption of dynamic neural fields to model curiosity, habituation, and persistence allows the robot to demonstrate various learning trajectories depending on the object. In addition, the approach exhibits interesting properties with respect to the learning of similar goals and the continuous switch between exploration and exploitation.
@@ I Introduction
Developmental robotics takes inspiration from various fields such as developmental psychology, neuroscience, machine learning, or philosophy. A popular domain of interest in the field and one of the core ideas of this study is called Curiosity. Curiosity can be seen as a particular case of intrinsic motivation (IM), and researchers have demonstrated that IM could be an efficient drive for the autonomous learning of behavior [1]. Moreover, the acquisition of different behaviors follows a developmental trajectory [2, 3]. Nowadays, there are multiple ways of modeling curiosity, and one of them consists of training predictors, such as forward and inverse models to create learning progress based on the prediction error. The use of curiosity in robotics supports the generation and learning of goals that increase in complexity [4, 5]. Goals are typically selected and managed via a top-down approach, where a central system monitors and determines the selection and learning time of the goals. Despite demonstrating impressive results, this approach narrows the understanding of curiosity, as it involves the interaction between various cognitive processes.
@@
Along with curiosity, attention is a fundamental cognitive process. In fact, attention allows humans to focus on specific elements of our environment. However, attention is a broad term that gathers several specific mechanisms. For example, visual attention is the ability to maintain gaze on prominent stimuli and develops during the first months after birth [6], [7]. To help search for a specific visual stimulus, inhibition of return is a well-studied mechanism that naturally prevents a person from looking twice at the same location in a short time interval [8]. In addition, we can distinguish two types of visual attention: exogenous, driven by external stimuli, or endogenous and goal-driven [9], [10]. These two types of attention are often referred to as bottom-up for the former and top-down for the latter, and inhibition of return can occur during both, but whether there are shared neural pathways between them to occur is still subject to debate [11]. In the brain, arousal is mediated by the Locus Coeruleus-Norepinephrine system (LC-NE) and demonstrates different patterns of neural activity. According to adaptive gain theory [12], LC neurons exhibit two modes of activation: tonic and phasic. During tonic activation, the subject disengages from the current task and engages in an exploratory behavior. In contrast, phasic LC activation engages in exploitation and thus is driven by task-specific outcomes. IM has been shown to be involved in the tonic and phasic activation of dopamine [13], overlapping certain neural pathways with noradrenaline (norepinephrine) [14] and thus could interact directly with the LC-NE system. If curiosity and more especially its learning progress component is related to attention, then the LC-NE system is a potential starting point.
@@
Here, we propose a robotic cognitive architecture for the online discovery and learning of goals, where a robotic arm is learning how to interact with simple objects. To do so, we identify and model several cognitive mechanisms to determine if the robot should explore and discover potential goals through bottom-up attention, or exploit and learn to achieve these goals. We introduce a component inspired by the LC-NE system to dynamically switch between goal discovery and learning. During both exploration and exploitation, an inhibition of return mechanism generates a set of actions toward the object location until producing a change in the environment. If the inhibition of return is shared by exploration and exploitation, there is, however, a conceptual difference between them since the discovery of a new goal in the environment rests on exogenous attention, and on a top-down and goal-driven approach for the learning of these goals. For the exploration stage, we complete the attentional mechanism connected to the LC-NE system with a habituation paradigm [15], where the discovery of a new goal produces a tonic activation and limits the exploration time. To balance this process, the absence of a new goal signals a shift from tonic to phasic activation. During the latter, the learning of a goal occurs with the use of simple neural networks that implement a forward and inverse model. The output of these neural networks determines an error associated with a goal and its learning progress. Except for the forward and inverse models, the entire architecture is designed with dynamic neural fields [16]. The goals are then represented within the neural fields by the neural activation of their learning progress and are projected to the LC-NE system, which then chooses to exploit the goal with the highest learning progress. By design, the model follows an enactive approach by continuously pushing the robot to interact with its environment, thus shaping its cognition using several forms of working memory. We will demonstrate that when similar goals are discovered, the robot autonomously only focuses on the one with the highest error and slowly forgets about the other. We evaluated several neural parameters such as persistence, prediction error inhibition, and habituation that modulate and optimize the learning of goals, depending on the objects. To our knowledge, this is the first attempt to model intrinsic motivation with dynamic neural fields and thus support the robot to express different learning behaviors by tuning intrinsic properties of the dynamic neural fields.
@@
This paper will provide a theoretical background according to the literature in Section II before presenting the architecture of the experiments (Section III). Then, we will present the experimental setup with the frameworks used in Section IV followed by the different results (Section V). Finally, we will discuss the advantages and limitations of the architecture (Section VI) before concluding.
@@ II-A Curiosity and Intrinsic motivation
In terms of biological perspective, intrinsic motivation (IM) should be distinguished from extrinsic motivation [17]. The former can be seen as a way for the brain to monitor the learning efficiency of a particular knowledge, while the latter directly guides the learning itself through external cues/rewards. Then, there are different metrics to compute intrinsic motivation [18], separating knowledge-based systems from competence-based systems [19]. In the literature, there is a consensus that classifies curiosity as a particular case of intrinsic motivation [20]. In fact, curiosity focuses on the learning progress hypothesis [21], [1], and formulates it as the evolution of prediction error while learning a task. In practice, curiosity can be used to learn a world model by focusing on the sensorimotor experience [22]. The terms goal and skill often refer to the same notion, even if a skill can be formed by the composition of several goals. For example, the grasping skill for a robot arm is formed by several goals, such as reaching, opening the gripper, adjusting the end-effector position, and closing the gripper. GRAIL [23] is a cognitive architecture operating at different levels where the IM used is competency-based, determining the best goal to pursue depending on how well the learning occurs. Intrinsic motivation in robotics is also used to form hypotheses in developmental psychology. For example, a developmental robotics model identified IM to explain the sudden improvement of tool-use in children [24]. However, the discovery of goals in visually rich environments is problematic as the exploration space rapidly expands, and an intrinsically motivated system loses efficiency. To address this issue, researchers are applying a set of dimension reduction methods. One of the first studies to overcome these limitations is the case of a robot arm interacting with a ball [25]. In this case, a variational autoencoder is trained, and the goals are sampled from the resulting goal space. Other works in the domain and related to robot grasping extended this concept by linking and compressing the object features, the action, and the sensori outcomes together [26]. By doing so, the robot creates a latent space where the learning of forward and inverse models is facilitated to achieve grasping. Working with high-dimensional spaces and intrinsic motivation is far from trivial, and the sampling problem can gain insight from a more developmental point of view. In these studies, it is necessary to perform several rounds of exploration to build up the different learning progress of goals and then select them for learning. However, it is not clear how the switch between discovery and learning takes place in childhood. In our work, we propose that the prediction errors contribute to some extent to easing this shift. Intrinsic motivation is a reproducing force of developmental patterns [27], but actively and efficiently sampling the environment remains difficult to tackle [28]. Finally, the top-down approach to select and control the learning of goals creates the homunculus problem [29]. In cognitive science, this principle is known to limit the understanding of a phenomenon, especially how the cognitive control of goals might be achieved. To address this issue, our research work search to model curiosity as a set of interconnected executive functions [30, 31] to provide flexibility for the discovery and learning of goals, and to apprehend the underlying question of cognitive control. As a starting point, researchers demonstrated that attention is involved in IM, and provides a perspective on a possible neural basis of intrinsic motivation [32].
@@ II-B Attention and inhibition of return
Visual attention is essential for an individual to focus on an unexpected event and to optimize cognitive processes on a specific task [10]. This distinction separates exogenous (bottom-up) from endogenous (top-down) attention. If motivation can sharpen exogenous spatial attention under certain conditions [33], the modulation of top-down attention can enhance the learning of a particular task [34, 35, 36]. The neural basis for these distinct processes is, for a significant part, shared [37], but not entirely [38]. Regarding the coupling between attention and intrinsic motivation, a study used reinforcement learning to allocate and shift attention [39]. Closer to our approach, researchers used bottom-up attention to learn object affordances and decompose a task into sub-goals [40]. Here, we choose to focus on the inhibition of return effect (IOR) to facilitate attention toward an object and ease the discovery of new goals. In the literature, the general mechanism of exploration is often summarized as random and direct exploration [41, 42] where the former consists of producing random actions with a high uncertainty but potentially large expected reward. The latter directs exploration toward more certain rewards but necessitates prior information. With curiosity in robotics, this difference can be translated to a motor babbling behavior for random exploration and the building of representations for direct exploration. For example, a robot can learn a representation of a diverse set of goals encountered during random exploration to generate new goals [26]. Here, our purpose is to propose an additional exploration method that can reduce the uncertainty of random exploration via the inhibition of return effect. This mechanism was pointed out by Posner [8] in visual attention when he discovered that the return of attention to a cue previously attended expresses a longer time. On the contrary, there is a facilitation (faster response time) if the time interval between cue-targets is short. This consequence endows the brain with a natural foraging mechanism of visual stimuli [43], but it remains necessary to proceed to an attentional disengagement to observe the IOR effect [44]. However, it appears that an endogenous disengagement of attention is not enough in order to observe the effect and is instead always relying on exogenous disengagement of attention [45]. In addition, it is possible to induce an inhibitory of return for both exogenous (uncued location) [46], [11] and endogenous (cued location) attention, but the effect seems highly correlated to the cue’s saliency and thus relies on an important exogenous component [47]. Regarding IOR in action production, there is no consensus on whether the effect is directly or partially involved. Several research reproduced the reaction time of IOR during a hand reaching movement [48] and even showed that reaching paths were biased toward the cue locations under certain conditions [49]. Directly connected to our research, it has been determined that IOR effects can be applied to objects [50]. Even more, inhibition can interact with working memory to mediate goal-directed action toward an object, optimizing performance for future interaction. In consequence, we intend to generate and use IOR around the object location so the robot can generate goal-driven actions. To be able to identify these two forms of attention and decide to engage in exploration, we take inspiration from the Locus Coeruleus.
II-C Exploration and exploitation
@@
The locus coeruleus-norepinephrine system (LC-NE) is a brainstem nucleus that is widely connected to the neocortex. The first unifying approach with respect to its role and functions is the adaptive gain theory [12]. The theory demonstrates two different neural activations, depending on the arousal state and the relation with task engagement. More precisely, the LC-NE system exhibits a tonic neural activation during exploration, raising the level of arousal and thus bottom-up attention to salient events [51]. Here, the exploration strategy consists of generating goal poses around the object with IOR and observing the possible outcomes. Phasic activation suggests a goal-oriented behavior (exploitation) with a focus on the task (top-down attention). Then, the LC-NE system is a crossroad between attentional processes and decision making. Aston-Jones and Cohen determine the emergence of a tonic activation when the utility of a task vanishes. The LC-NE system receives direct connection from the anterior cingulate (Acc) and the orbitofrontal cortices (OFC), which are directly involved in task-related utility and decision making [52], [53]. More specifically, the ACC appears to represent predictions of value about reward and, more generally, uncertainty [54]. There is a consensus on the recognition of the basal ganglia as the reward processing site that can support reinforcement learning [55]. Several robotics experiments take inspiration from the phasic delivery of dopamine to signal rewards based on prediction errors [56], [57]. We can observe a similarity between the phasic delivery of noradrenaline from the LC and the phasic dopamine-based signal emerging from the basal ganglia, especially after pointing out that dopamine and noradrenaline share neural pathways [58]. In addition, researchers demonstrate that if dopamine and noradrenaline are related to reward prediction, only the latter is predictive of task engagement and is strongly activated when cues indicate a new task condition. This means that, contrary to dopamine, noradrenaline is strongly involved in motivating current or future engagement. The Acc directly projects prediction information about the task utility to the LC-NE system, which in turn decides if the utility is high enough to drive the neural activation in a phasic mode, hence to an exploitatory stage. Researchers proposed an alternative view of the LC by describing its phasic activation as a reset network capable of reorienting resources to a new task [59, 51] while still being strongly sensitive to rewards [60, 61]. Thus, integrating the Aston-Jones LC system theory with prediction-error sensitivity appears promising for enhancing curiosity-driven robotics. It will allow the robot to engage in exploitation depending on uncertainty, meaning on how accurate the robot’s predictions are, based on the learning progress. As a consequence, the LC model will follow a phasic neural activation, thus autonomously tuning attention to modulate the level of arousal and engage in the learning of a goal. Recent neuroscience research highlights the role of the median raphe nucleus (MRN), a brainstem nucleus, in perseverative, exploratory, and disengaged states [62]. In their study, they discovered that GABAergic and serotonergic neurons influence perseverative states and exploration, respectively. There are currently no studies that investigate the influence of the MRN on the LC-NE system and vice versa. Their close proximity in the brain and their implication in exploratory and exploitative states could indicate a relationship, but this hypothesis is currently a speculation. We intend to use these findings to modulate exploration and regulate goal learning by investigating the role of excitatory and inhibitory neurons, respectively. To do so, we first identify habituation as a cognitive mechanism that can regulate exploration. Then, we investigate the process of persistence (i.e., perseverance) to regulate the allocated time for the learning of a goal.